# Improving threshold for fault-tolerant color code quantum computing by flagged weight optimization

### v1

@@ Abstract
Color codes are promising quantum error correction (QEC) codes because they have an advantage over surface codes in that all Clifford gates can be implemented transversally. However, thresholds of color codes under circuit-level noise are relatively low mainly because measurements of their high-weight stabilizer generators cause an increase in a circuit depth, and thus, substantial errors are introduced. This makes color codes not the best candidate. Here, we propose a method to suppress the impact of such errors by optimizing weights of decoders using flag qubits and reducing the circuit depth using cat states. We set the weights based on conditional error probabilities conditioned on the measurement outcomes of flag qubits. In numerical simulations, we improve the threshold of the (4.8.8) color code under the circuit-level noise from 0.14% to around 0.27%, which is calculated by using an integer programming decoder. Furthermore, in the (6.6.6) color code, we achieved a circuit-level threshold of around 0.36%, which is almost the same value as the highest value in the previous studies employing the same noise model. In both cases, the achieved logical error rates at low physical error rates are almost one order of magnitude lower than a conventional method that uses a single ancilla qubit for each stabilizer measurement. This method can also be applied to other weight-based decoders, making the color codes more promising for the candidate of experimental implementation of QEC. Furthermore, one can utilize this approach to improve a threshold of wider classes of QEC codes, such as high-rate quantum low-density parity check codes.
@@ I Introduction
Quantum computers have the potential to efficiently solve computationally difficult problems, such as factorization of large numbers [1] and simulations of quantum many-body systems [2]. However, qubits are highly susceptible to noise, making it difficult to perform accurate quantum computations. Quantum error correction (QEC) is a critical solution to suppress the impact of such a noise, enabling fault-tolerant quantum computing (FTQC) by encoding fragile physical qubits into robust logical qubits through quantum error correction codes (QECCs). In the theory of QEC, if an error probability per quantum gate is below a certain threshold, we can perform arbitrarily accurate quantum computations by increasing the number of physical qubits. Consequently, extensive research has been undertaken to establish FTQC protocols with a high threshold.
Currently, surface codes [3] are considered to be one of the most promising QECCs, as have been experimentally demonstrated with small code distances in recent years [4, 5, 6, 7]. The notable advantages of surface codes are their ease of physical implementation as well as their high thresholds. The thresholds of surface codes under the circuit-level noise are estimated to be around 0.5%-1.1% [8, 9, 10, 11, 12, 13], depending on each noise assumption and way of performing QEC. On the other hand, surface codes also have a drawback in terms of fault-tolerant implementation of logical gates. In order to realize large-scale FTQC, it is needed to implement a universal set of logical gates fault-tolerantly with low spatial and temporal overheads. However, even for certain Clifford gates, fault-tolerantly implementing logical gates using surface codes requires costly techniques [14, 15, 16], which can lead to significant overheads.
Another promising QECC is color codes [17, 18], which admit transversal implementation of all logical Clifford gates due to their high symmetry of stabilizer operators [19]. This property has led to color codes being considered as promising QECCs for achieving FTQC with low overhead. However, low thresholds of color codes have made the practical implementation of color code-based FTQC difficult. For two typical color codes, the (4.8.8) color code and the (6.6.6) color code, the thresholds under the circuit-level noise are around 0.08%-0.14% [20, 21] and 0.2%-0.47% [10, 22, 23, 24], respectively. The main cause of the low thresholds is that stabilizer generators of color codes are high-weight; in other words, they act on many data qubits. High-weight stabilizer generators cause an increase in the circuit depth of a syndrome measurement circuit, and thus, substantial errors are introduced.
@@
In this paper, we propose a method to improve thresholds of color codes under the circuit-level noise by optimizing the weights of a decoder using flag qubits [28, 29, 30, 31]. A flag qubit is an additional ancilla qubit that provides information about errors occurring on ancilla qubits, which allows us to correct more errors in the subsequent QEC procedure. We set weights for data errors and measurement errors based on conditional error probabilities conditioned on the measurement outcomes of local flag qubits. The conditional error probabilities are estimated by repeatedly executing a tailored quantum circuit offline prior to the decoding and obtaining information about errors from its measurement outcomes. This estimation method is efficient, accurate, and can be used even when the underlying noise is a priori unknown. We also perform a deflagging procedure [32, 33] to further improve the performance. Furthermore, our way of measuring syndrome using cat states reduces the circuit depth of a syndrome measurement circuit, and thus, suppresses the occurrence of errors. At the same time, the number of data qubits interacting with an ancilla qubit decreases, thus the propagation of errors is suppressed. As a decoder, we use the integer programming decoder that is not efficient throughout this paper. This is because our focus is not on decoders themselves but on how their performance improves by the proposed method. Although our method can be applied to other weight-based decoders, the simplicity of the formulation in the integer programming decoder makes the proposed method more straightforward to understand. Our approach is versatile in that it can also be applied to other QECCs that have high-weight stabilizer generators, such as high-rate quantum low-density parity check (LDPC) codes [34].
@@
In numerical simulations, we improve the threshold of the (4.8.8) color code under the circuit-level noise from 0.14% to around 0.27%, which is calculated by using the integer programming decoder. In addition, in the (6.6.6) color code, we achieved the circuit-level threshold of around 0.36%, which is almost the same as the highest value in the previous studies that employ the same noise model. Moreover, in both cases, the achieved logical error rates at low physical error rates are almost one order of magnitude lower than the conventional method that uses a single ancilla qubit for each stabilizer measurement and employs uniform weights.
@@
The rest of the paper is organized as follows. In Sec. II, we first introduce two types of 2D color codes. Then, we describe the decoding formulation and conventional weights of a decoder under the circuit-level noise using the integer programming decoder. A conventional syndrome measurement circuit is also shown. In Sec. III, we describe the details of the procedure for the proposed flagged weight optimization. Additionally, we explain the deflagging procedure and how to estimate the conditional error probabilities required for setting the weights. In Sec. IV, we show the numerical results for the logical error rates achieved by our method and discuss comparisons with existing methods. Finally, a conclusion is presented in Sec. V.
II Preliminary
@@ II.1 Color codes
$G_{f}^{X}:=\prod_{v\in f}X_{v},$ (1)
$G_{f}^{Z}:=\prod_{v\in f}Z_{v}.$ (2)
The code space is the simultaneous $+1$ eigenstate of all stabilizer operators. There are two main types of lattices that can be used to define 2D color codes: (4.8.8) lattice and (6.6.6) lattice. The (4.8.8) lattice is a semi-regular lattice where each vertex is incident to a square and two octagonal faces, and the (6.6.6) lattice is a regular lattice where each vertex is incident to three hexagonal faces. The 2D color codes defined on the (4.8.8) lattice and the (6.6.6) lattice with boundaries, in which a single logical qubit is encoded, are shown in Figs. 1 and 2, respectively. In these 2D color codes, we can implement all logical Clifford gates transversally [19].
@@ II.2 Decoding under circuit-level noise using integer programming decoder
×$\times\sum_{\begin{subarray}{c}P_{1},P_{2}\in\{I,X,Y,Z\}\\ P_{1}\otimes P_{2}\neq I\otimes I\end{subarray}}(P_{1}\otimes P_{2})\rho_{2}(P_{1}\otimes P_{2}),$ (7)
@@
    Each state preparation fails with the probability $p$, and an orthogonal state is prepared.
@@
    Each measurement fails with the probability $p$, and the measurement outcome is inverted.
@@
The task of the decoding is to estimate the most likely errors given a syndrome. Under the circuit-level noise, a syndrome is not reliable due to the presence of measurement errors. Thus, if a single syndrome measurement is performed and the decoding is carried out based on the syndrome, the decoding accuracy significantly gets worse. In this situation, it is possible to suppress performance deterioration due to measurement errors by repeating the syndrome measurement multiple times (in this study, $d$ times). Then, we decode $X$ and $Z$ errors separately based on the obtained syndrome in spacetime. In the case of the noise model where probabilities of data errors and measurement errors occurring are assumed to be independently and identically distributed (i.i.d.), i.e., phenomenological noise model, we can estimate the most likely errors by minimizing the total number of errors given the syndrome. However, under the circuit-level noise, the probabilities of data errors and measurement errors occurring are not identical because the way errors occur and propagate differs for each qubit. Also, they are not independent due to the correlated errors caused by entangling gates. Therefore, to identify the most likely errors, it is necessary to estimate the errors based on an actual error probability distribution.
@@
In the following, we explain how to decode under the circuit-level noise using the integer programming decoder, where the decoding problem is formulated as an integer programming problem. The decoding formulation is an extension of that in the phenomenological noise model provided in Ref. [35]. Here, we explain only the decoding of $X$ errors, but the decoding of $Z$ errors is also performed using the same algorithm. In this decoder, errors and syndrome values are represented as binary variables. Here, we define the binary variable $u^{(t)}_{v}\in\{0,1\}$ as the cumulative data error that has occurred in the qubit at the vertex $v$ until time $t$, and the binary variable $r^{(t)}_{f}\in\{0,1\}$ as the measurement error occurred on the face $f$ at time $t$. Then, we can express the measured syndrome value on the face $f$ at time $t$ as the binary value $s^{(t)}_{f}\in\{0,1\}$:
$s^{(t)}_{f}=\bigoplus_{v\in f}u^{(t)}_{v}\oplus r^{(t)}_{f}.$ (8)
@@
The syndrome difference between time $t$ and time $t-1$, i.e., $s^{(t)}_{f}\oplus s^{(t-1)}_{f}$ detects the data errors occured at time $t$. Thus, the equation that the syndrome values should satisfy is given as follows:
$\bigoplus_{v\in f}x^{(t)}_{v}\oplus r^{(t)}_{f}\oplus r^{(t-1)}_{f}=s^{(t)}_{f}\oplus s^{(t-1)}_{f}\quad\forall f,$ (9)
@@
where the binary variable $x^{(t)}_{v}\in\{0,1\}$ denotes newly occuring data error at time $t$, which corresponds to the XOR of $u^{(t)}_{v}$ and $u^{(t-1)}_{v}$. By setting weights for each error and minimizing the weighted number of errors satisfying Eq. (9), it is possible to decode based on an error probability distribution. Thus, the decoding problem under the circuit-level noise is formulated as an integer programming problem as follows:
@@
$\mathrm{s.t.}\quad\bigoplus_{v\in f}x^{(t)}_{v}\oplus r^{(t)}_{f}\oplus r^{(t-1)}_{f}=s^{(t)}_{f}\oplus s^{(t-1)}_{f}\quad\forall f.$ (11)
@@
Note that an error probability distribution used in Eqs. (12) and (13) needs to be estimated. So far, several methods have been proposed to estimate the error probability distribution in this context. Refs.  [13, 33, 36, 24] have proposed methods to estimate the error probability distribution by analytically counting possible error events. In Ref. [37, 38], by calculating the expected values of the syndrome from sufficient syndrome data, the error probability distribution is estimated. In Ref. [39], the error probability distribution is estimated by repeatedly decoding the obtained syndrome.
@@ II.3 Conventional syndrome measurement circuit
Figure 4: CNOT order for the (4.8.8) color code with a single ancilla qubit for each face. The black numbers indicate the time steps in which the CNOT gates are applied. The black lines represent interactions between qubits. The total number of qubits required for this method is $(3d^{2}+6d-5)/4$.
Figure 5: CNOT order for the (6.6.6) color code with a single ancilla qubit for each face. The total number of qubits required for this method is $(9d^{2}-1)/8$.
Syndrome measurements need to be performed without destroying the encoded state. If data qubits are directly measured, the superposition state will be destroyed, so indirect measurements are used for syndrome measurements. In the conventional method [9, 20], measurements of $X$ and $Z$ stabilizer generators are performed using a single ancilla qubit, as shown in Figs. 3(a) and 3(b), respectively. In this method, the more data qubits the stabilizer generators act on, the deeper the circuit depth of the syndrome measurement circuit, resulting in more locations where errors may occur. Also, the more data qubits interact with an ancilla qubit, the more error propagation paths there are, so that an error can propagate widely. Thus, color codes with high-weight stabilizer generators are more susceptible to errors compared to surface codes, which have at most weight-four stabilizer generators. This is the reason why the thresholds of color codes are low under the circuit-level noise. In Ref. [20], it has been shown that the threshold of the (4.8.8) color code using this syndrome measurement circuit is around 0.08%. Examples of the CNOT order are shown for the (4.8.8) color code in Fig. 4 and for the (6.6.6) color code in Fig. 5. We also show the $X$ stabilizer measurement circuits for each face of the (4.8.8) color code in Fig. 6 and of the (6.6.6) color code in Fig. 7. $Z$ stabilizer measurement circuits are similar, except that the basis changes to the $Z$-basis, and the direction of the CNOT gates is reversed. We need to implement eight depth of the CNOT gates in the case of the (4.8.8) color code and six depth of the CNOT gates in the case of the (6.6.6) color code.

### v2

@@ Abstract
Color codes are promising quantum error correction (QEC) codes because they have an advantage over surface codes in that all Clifford gates can be implemented transversally. However, thresholds of color codes under circuit-level noise are relatively low mainly because measurements of their high-weight stabilizer generators cause an increase in a circuit depth, and thus, substantial errors are introduced. This makes color codes not the best candidate for fault-tolerant quantum computing. Here, we propose a method to suppress the impact of such errors by optimizing weights of decoders using conditional error probabilities conditioned on the measurement outcomes of flag qubits. In numerical simulations, we improve the threshold of the (4.8.8) color code under circuit-level noise from 0.14% to around 0.27%, which is calculated by using an integer programming decoder. Furthermore, in the (6.6.6) color code, we achieve a circuit-level threshold of around 0.36%, which is almost the same value as the highest value in the previous studies employing the same noise model. In both cases, an effective code distance is also improved compared to a conventional method that uses a single ancilla qubit for each stabilizer measurement. Thereby, the achieved logical error rates at low physical error rates are almost one order of magnitude lower than those of the conventional method with the same code distance. Even when compared to the single ancilla method with higher code distance, considering the increased number of qubits used in our method, we achieve lower logical error rates in most cases. This method can also be applied to other weight-based decoders, making the color codes more promising for the candidate of experimental implementation of QEC. Furthermore, one can utilize this approach to improve a threshold of wider classes of QEC codes, such as high-rate quantum low-density parity check codes.
@@ I Introduction
Quantum computers have the potential to efficiently solve computationally difficult problems, such as factorization of large numbers [1] and simulations of quantum many-body systems [2]. However, qubits are highly susceptible to noise, making it difficult to perform accurate quantum computations. Quantum error correction (QEC) is a critical solution to suppress the impact of noise, enabling fault-tolerant quantum computing (FTQC) by encoding fragile physical qubits into robust logical qubits through quantum error correction codes (QECCs). In the theory of QEC, if an error probability per quantum gate is below a certain threshold, we can perform arbitrarily accurate quantum computations by increasing the number of physical qubits. Consequently, extensive research has been undertaken to establish FTQC protocols with a high threshold.
Currently, surface codes [3] are considered to be one of the most promising QECCs, as they have been experimentally demonstrated in recent years [4, 5, 6, 7]. The notable advantages of surface codes are their ease of physical implementation as well as their high thresholds. The thresholds of surface codes under circuit-level noise are estimated to be around 0.5%-1.1% [8, 9, 10, 11, 12, 13], depending on each noise assumption and way of performing QEC. On the other hand, surface codes also have a drawback in terms of fault-tolerant implementation of logical gates. In order to realize large-scale FTQC, it is needed to implement a universal set of logical gates fault-tolerantly with low spatial and temporal overheads. However, even for certain Clifford gates, fault-tolerantly implementing logical gates using surface codes requires costly techniques [14, 15, 16], which can lead to significant overheads.
Another promising QECC is color codes [17, 18, 19], which admit transversal implementation of all logical Clifford gates due to their high symmetry of stabilizer operators [20]. This property has led to color codes being considered as promising QECCs for achieving FTQC with low overhead. Experimentally, color codes with small code distances have been demonstrated in recent years [21, 22, 7]. However, low thresholds of color codes have made the practical implementation of color code-based FTQC difficult. For two typical color codes, the (4.8.8) color code and the (6.6.6) color code, the thresholds under circuit-level noise are around 0.08%-0.14% [23, 24] and 0.2%-0.47% [10, 25, 26, 27, 28], respectively. In the [[4,2,2]]-concatenated toric code, which is a subsystem version of the (4.8.8) color code, a circuit-level threshold of 0.41% has been obtained [29] due to the utilization of the gauge degree of freedom, but it does not support transversal $H$ and $S$ gate. The main cause of the low thresholds in color codes is that stabilizer generators of color codes are high-weight; in other words, they act on many data qubits. High-weight stabilizer generators cause an increase in the circuit depth of a syndrome measurement circuit, and thus, substantial errors are introduced.
@@
In this paper, we propose flagged weight optimization (FWO), a method to improve thresholds of color codes under circuit-level noise by optimizing the weights of a decoder using conditional error probabilities conditioned on the measurement outcomes of flag qubits. A flag qubit [33, 34, 35, 36, 37] is an additional ancilla qubit that provides information about errors occurring on ancilla qubits, which allows us to correct more errors in the subsequent QEC procedure. We set weights for data errors and measurement errors based on conditional error probabilities conditioned on the measurement outcomes of local flag qubits. Ref. [10] also uses flag information in the decoding graph, but there are a lot of differences from our method. In Ref. [10], flag edges are added and each weight is set by a conventional way, except for a renormalization for the weights of edges that are unlikely to flip, whereas we do not add additional nodes and edges. We also propose a method to estimate the conditional error probabilities, which is efficient, accurate, and can be used even when the underlying noise is a priori unknown. The conditional error probabilities are estimated by repeatedly executing a tailored quantum circuit offline prior to the decoding and obtaining information about errors from its measurement outcomes. In addition, we perform deflagging procedure proposed in Refs. [38, 39] combined with FWO to further improve the performance, which is a new application of deflagging.
In our syndrome measurement circuits using flag qubits, cat states are prepared first. The use of cat states reduces the circuit depth of syndrome circuits, and thus suppresses the occurrence of errors, as also mentioned in Ref. [24]. At the same time, the number of data qubits interacting with an ancilla qubit decreases, thus the propagation of errors is suppressed. As to the implementation on hardware, cat states reduce the connectivity requirement, which is especially essential in superconducting architectures. As a decoder, we use the integer programming decoder, which is not efficient, throughout this paper. This is because our focus is not on decoders themselves but on how their performance improves by the proposed method. Although our method can be applied to other weight-based decoders, the simplicity of the formulation in the integer programming decoder makes the proposed method more straightforward to understand. Our approach is versatile in that it can also be applied to other QECCs that have high-weight stabilizer generators, such as high-rate quantum low-density parity check (LDPC) codes [40].
@@
In numerical simulations, we assess logical error rates for memory errors and do not analyze time-like logical errors rates [41], which may be relevant for implementing non-Clifford gates. We improve the threshold of the (4.8.8) color code under circuit-level noise from 0.14% to around 0.27%, which is calculated by using the integer programming decoder. In addition, in the (6.6.6) color code, we achieved the circuit-level threshold of around 0.36%, which is the same within statistical errors as the highest value of 0.37(1)[26], among the previous studies that employ the same noise model. Note that our threshold values are calculated by using code distances up to $d=9$. Moreover, in both cases, an effective code distance is also improved compared to a conventional method that uses a single ancilla qubit for each stabilizer measurement and employs uniform weights, meaning that FWO helps correct large hook errors that arise from relatively few faults. Thereby, the achieved logical error rates at low physical error rates are almost one order of magnitude lower than those of the conventional method when comparing with the same code distance. We also verified that our method achieves lower logical error rates in most cases, even when comparing to the conventional method with a higher code distance, assuming a given number of available qubits where the conventional method with a higher code distance is feasible.
@@
The rest of the paper is organized as follows. In Sec. II, we first introduce 2D color codes. Then, we describe the decoding formulation and conventional weights of a decoder under circuit-level noise using the integer programming decoder. We also explain conventional syndrome measurement circuits for color codes and the reasons why they lead to a low threshold. In Sec. III, we describe the details of the procedure for the proposed FWO. Additionally, we explain the deflagging procedure and how to estimate the conditional error probabilities required for setting the weights. In Sec. IV, we show the numerical results for the logical error rates achieved by our method and discuss comparisons with existing methods. Finally, a conclusion is presented in Sec. V.
II Preliminaries
@@ II.1 Color codes
$G_{f}^{X}:=\prod_{v\in\delta f}X_{v},$ (1)
$G_{f}^{Z}:=\prod_{v\in\delta f}Z_{v}.$ (2)
Here, $\delta f$ is the set of vertices that touch $f$. The code state is the simultaneous $+1$ eigenstate of all stabilizer operators. There are three types of lattices that can be used to define 2D color codes: (4.8.8) lattice, (6.6.6) lattice, and (4.6.12) lattice [23]. Among these lattices, we focus on (4.8.8) lattice and (6.6.6) lattice in this work. The (4.8.8) lattice is a semi-regular lattice where each vertex is incident to a square and two octagonal faces, and the (6.6.6) lattice is a regular lattice where each vertex is incident to three hexagonal faces. The 2D color codes defined on the (4.8.8) lattice and the (6.6.6) lattice with boundaries, in which a single logical qubit is encoded, are shown in Figs. 1 and 2, respectively. In these 2D color codes, we can implement all logical Clifford gates transversally [20].
@@ II.2 Decoding under circuit-level noise using integer programming decoder
×$\times\sum_{\begin{subarray}{c}P_{1},P_{2}\in\{I,X,Y,Z\}\\ P_{1}\otimes P_{2}\neq I\otimes I\end{subarray}}(P_{1}\otimes P_{2})\rho_{2}(P_{1}\otimes P_{2}),$
@@
    Each state preparation fails with probability $p$, and an orthogonal state is prepared.
@@
    Each measurement fails with probability $p$, and the measurement outcome is inverted.
@@
The task of a decoding is to estimate the most likely errors given a syndrome. Under the circuit-level noise, a syndrome is not reliable due to the presence of measurement errors. Thus, if a single syndrome measurement is performed and the decoding is carried out based on the syndrome, the decoding accuracy significantly gets worse. In this situation, it is possible to suppress performance deterioration due to measurement errors by repeating the syndrome measurement multiple times (in this study, $d$ times). Then, we decode $X$ and $Z$ errors separately based on the obtained syndrome in spacetime. In the case of a noise model where probabilities of data errors and measurement errors occurring are assumed to be independently and identically distributed (i.i.d.), i.e., phenomenological noise model, we can estimate the most likely errors by minimizing the total number of errors given the syndrome. However, under the circuit-level noise, the probabilities of cumulative data errors and measurement errors for each round, i.e., the probabilities that edges in the decoding graph are triggered, are not identical because the way errors occur and propagate differs for each qubit. Also, they are not independent due to the correlated errors caused by entangling gates. Therefore, to identify the most likely errors, it is necessary to estimate the errors based on an actual error probability distribution.
@@
In the following, we explain how to decode under the circuit-level noise using the integer programming decoder, where the decoding problem is formulated as an integer programming problem. The decoding formulation is an extension of that in the phenomenological noise model provided in Ref. [42]. Here, we explain only the decoding of $X$ errors, but the decoding of $Z$ errors is also performed using the same algorithm. In this decoder, errors and syndrome values are represented as binary variables. Here, we define the binary variable $u^{(t)}_{v}\in\{0,1\}$ as the cumulative data error that has occurred on the qubit at vertex $v$ until time $t$, and the binary variable $r^{(t)}_{f}\in\{0,1\}$ as the measurement error occurred on the face $f$ at time $t$. Then, we can express the measured syndrome value on the face $f$ at time $t$ as the binary value $s^{(t)}_{f}\in\{0,1\}$:
$s^{(t)}_{f}=\bigoplus_{v\in\delta f}u^{(t)}_{v}\oplus r^{(t)}_{f}.$ (5)
@@
The syndrome difference between time $t$ and time $t-1$, i.e., $s^{(t)}_{f}\oplus s^{(t-1)}_{f}$ detects the data errors occurred at time $t$. Thus, the equation that the syndrome values should satisfy is given as follows:
$\bigoplus_{v\in\delta f}x^{(t)}_{v}\oplus r^{(t)}_{f}\oplus r^{(t-1)}_{f}=s^{(t)}_{f}\oplus s^{(t-1)}_{f}\quad\forall f,$ (6)
@@
where the binary variable $x^{(t)}_{v}\in\{0,1\}$ denotes newly occuring data error at time $t$, which corresponds to the XOR of $u^{(t)}_{v}$ and $u^{(t-1)}_{v}$. If $x^{(t)}_{v}=1$, it indicates that a data error has newly occurred on the corresponding data qubit at time $t$. Conversely, if $x^{(t)}_{v}=0$, it indicates that a data error has not newly occurred on the corresponding data qubit at time $t$. By setting weights for each error and minimizing the weighted number of errors satisfying Eq. (6), it is possible to decode based on an error probability distribution. Thus, the decoding problem under the circuit-level noise is formulated as an integer programming problem as follows:
@@
$\mathrm{s.t.}\quad\bigoplus_{v\in\delta f}x^{(t)}_{v}\oplus r^{(t)}_{f}\oplus r^{(t-1)}_{f}=s^{(t)}_{f}\oplus s^{(t-1)}_{f}\quad\forall f.$ (8)
@@
Note that an error probability distribution used in Eqs. (9) and (10) needs to be estimated. So far, several methods have been proposed to estimate the error probability distribution in this context. Refs.  [13, 39, 43, 28] have proposed methods to estimate the error probability distribution by analytically counting possible error events. In Refs. [44, 45], the error probability distribution is estimated by calculating the expected values of the syndrome from appropriate syndrome data. In Ref. [46], the error probability distribution is estimated by repeatedly decoding the obtained syndrome.
@@ II.3 Conventional syndrome measurement circuit
Figure 4: CNOT order for the (4.8.8) color code with a single ancilla qubit for each face. The black numbers indicate the time steps in which the CNOT gates are applied. The black lines represent interactions between qubits. The total number of qubits required for this method is $n_{4.8.8}^{\text{single}}(d)=(3d^{2}+6d-5)/4$.
Figure 5: CNOT order for the (6.6.6) color code with a single ancilla qubit for each face. The total number of qubits required for this method is $n_{6.6.6}^{\text{single}}(d)=(9d^{2}-1)/8$.
Syndrome measurements need to be performed without destroying the encoded state. If data qubits are directly measured, the superposition state will be destroyed, so indirect measurements are used for syndrome measurements. In the conventional method [9, 23], measurements of $X$ and $Z$ stabilizer generators are performed using a single ancilla qubit, as shown in Figs. 3(a) and 3(b), respectively. In this method, the more data qubits the stabilizer generators act on, the deeper the circuit depth of the syndrome measurement circuit, resulting in more locations where errors may occur. Also, the more data qubits interact with an ancilla qubit, the more error propagation paths there are, so that an error can propagate widely. Here, an error propagation path is a set of locations in a circuit where an occurring error propagates. Thus, color codes with high-weight stabilizer generators are more susceptible to errors compared to surface codes, which have at most weight-four stabilizer generators. This is the reason why the thresholds of color codes are low under the circuit-level noise. Here, we consider a CNOT schedule where the $X$ stabilizer measurement is performed first, followed by the $Z$ stabilizer measurement. In Ref. [23], it has been shown that the threshold of the (4.8.8) color code using this syndrome measurement circuit is around 0.08%. Examples of a possible CNOT order are shown for the (4.8.8) color code in Fig. 4 and for the (6.6.6) color code in Fig. 5. The total number of required qubits, including both data and ancilla qubits, is $n_{4.8.8}^{\text{single}}(d)=(3d^{2}+6d-5)/4$ for the (4.8.8) color code and $n_{6.6.6}^{\text{single}}(d)=(9d^{2}-1)/8$ for the (6.6.6) color code. We also show the $X$ stabilizer measurement circuits for each face of the (4.8.8) color code in Fig. 6 and of the (6.6.6) color code in Fig. 7. $Z$ stabilizer measurement circuits are similar, except that the basis changes to the $Z$-basis, and the direction of the CNOT gates is reversed. We need to implement CNOT gates with a depth of 8 for the (4.8.8) color code and a depth of 6 for the (6.6.6) color code.